• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 16:04:59 on March 21, 2005

    Entry number 142116

    keyword=Beam Time Accounting
    
    	
    Hall A

    03/21/05 0800-1600
    Experiment #: e01015
    Shift Leader: J. P. Chen
    Shift Workers: X. Jiang, M. Potokar

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    0.130.000.000.000.000.080.250.500.97
    Beam Available
    but Not in Use
    0.000.000.000.000.000.250.670.000.92
    Beam Not Available
    or Unacceptable
    0.580.000.000.000.000.080.080.080.82
    Accelerator Con-
    figuration Change
    0.281.001.001.001.000.580.000.425.28
    Experiment
    Ready
    1.001.001.001.001.000.750.331.007.08
    Planned Config-
    uration Change
    0.000.000.000.000.000.000.000.000.00
    Unplanned
    Experiment Down
    0.000.000.000.000.000.250.670.000.92
    Hall not
    ready reasons
    Target0.000.000.000.000.000.000.000.000.00
    Data Acq0.000.000.000.000.000.000.080.000.08
    Other0.000.000.000.000.000.000.000.000.00
    HRSL0.000.000.000.000.000.000.000.000.00
    HRSR0.000.000.000.000.000.250.580.000.83
    HALL_OFF0.000.000.000.000.000.000.000.000.00

    Comments: 8:00 no beam, vacuum pump tripped. 8:35-8:43 data taking 8:43-13:25 beam study, hall controlled access to fix Septum problem. 13:25-13:35 asked for 5 uA, did Harp Scan 14:35 right dipole set, asked for 40 uA. 14:43 production 15:35 asked for beam not through Compton Chicane. 16:00 stiil beam tuning

    Data Entry by: jpchen